The table below describes the smoking habits of a group of asthma sufferers. If one of the 1156 people is randomly selected, find the probability that the person is a man or a heavy smoker. Round to three decimal places as needed Occasional Regular Heavy Smoker Smoker Smoker Total 601 Nonsmoker 431 382 813 Men Women Total 50 48 98 71 86 157 49 39 1156 OA, 0.557 O B. 0.596 O c. 0.511 O D. 0.554
Answer #1

Solution:- (D)

The probability that the person is a man or a heavy smoker is 0.554.

Total number of people = 1156

Number of person who are man or smokers = 601 + 88 - 49 = 640

The probability that the person is a man or a heavy smoker = 640/1156 = 0.554